OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (-3, -1) with a slope of 4.

OpenStudy (anonymous):

y = 4x - 11 y = 4x + 11 y = 4x + 7 y = 4x - 7

OpenStudy (anonymous):

\[ \text{ Substitute the given slope and point into the formula } y = mx + b\\\\ \text{ Then find the value of b. }\\\\ -1= 4(-3) + b b= 11\\\\ \text{ Then just substitute the slope In place of m, and the value of the y intercept In place of b }\\\\ \text{ In the equation y = mx + b }\\\\ y = x +11 \]

OpenStudy (anonymous):

you welcome !